A painter can finish painting a house in 4 hours. Her assistant takes 6 hours to finish the same job. How long would it take for them to complete the job if they were working together?
Answer by Ruli(21)

The painter = 1/4x
Her assistant = 1/6x
1/4x + 1/6x = 1
Multiply each side by the LCD, 12
12(1/4x) + 12(1/6x) = 12(1)
3x + 2x = 12
5x = 12
Divide each side by 5
x = 12/5
Working together, the painter and her assistant can paint the house in 12/5 hr.
check:
1/4(12/5) + 1/6(12/5) = 1
12/20 + 12/30 = 1
36/60 + 24/60 = 1
60/60 = 1
1 = 1